OpenStudy (anonymous):

Find critical points h(x)= root over(6-3x^2)

OpenStudy (zehanz):

To find critical points, you have to differentiate h. Then solve h'(x)=0. The solutions are the critical points.

OpenStudy (anonymous):

HI , i did and sort of stuck at the end -3x(6-3x^2)^-1/2=0, x=0, and (6-3x^2)^(-1/2)=0, so how do i do at the second part... thanks for the help.

OpenStudy (anonymous):

can you retype the function? as written, this is what you have: h(x)= root over(6-3x^2)

OpenStudy (zehanz):

OK, so you've got your derivative! Solving h'(x) = 0: \[-\frac{ 3x }{ \sqrt{6-3x^2} }=0 \Leftrightarrow 3x=0\] So x=0

OpenStudy (zehanz):

Remember: If a/b=0 this means: a=0 (and b is not 0). So there is an easy solution here...

OpenStudy (anonymous):

so, there is only one critical number (0)..

OpenStudy (zehanz):

Yes, we can get no more from this one...

OpenStudy (anonymous):

so what is critical point is sqrt 6 !

OpenStudy (anonymous):

(0, sqrt6)

OpenStudy (zehanz):

Here is the graph, it confirms what you found:

OpenStudy (anonymous):

its webwork solution and it is not getting right..

OpenStudy (zehanz):

Thanks!

OpenStudy (anonymous):

what would be the critical point ?

OpenStudy (zehanz):

(0, sqrt(6))

OpenStudy (zehanz):

There is even a maximum in that point.

OpenStudy (anonymous):

the answer (0, sqrt(6)) cameout wrong in the webwork..

OpenStudy (zehanz):

But the real question now is: are you convinced that it is true?